2. Plot these data as electromagnetic waves on a graph and find the following:


a. Period


b. Frequency



2.1 Time= 15s


Cycles= 5


2.2 Time= 30s


Cycles= 6s



3. The wavelength of an electromagnetic wave is 450nm, and the speed of light (c) is 3x108m/s. Determine the frequency, energy, and numbers of the electron.






1
Expert's answer
2022-05-03T01:53:03-0400

Period=Time/Cycles=15/5=3 s

Frequency=Cycles/Time=5/15=1/3



3."\nu=c/\lambda=3x10^8/(450x10^{-9})=6.67 x10^{14}s^{-1}"

"E=h\nu=6.63 x10^{-34}x6.67x10^{14}=44.22 x10^{-20}J"

Electron number "44.22x10^{-20}\/1.6 x10^{-19}=2.8" electrons.